OpenStudy (anonymous):

Crystal is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 12. To give everyone a good view, each row must have three more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 3(x - 1). Explain to Crystal how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the thirtieth row. Use complete sentences.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@phi

OpenStudy (phi):

This is like a puzzle. The first step is write down how many seats in, for example, the first 3 rows... we have to look for a pattern

OpenStudy (anonymous):

Hmmm. Can you help me out a little with that?

OpenStudy (phi):

write down the labels row # # of seats f(x) s(x)=3(x-1) under row # start with 1 row # # of seats f(x) s(x)=3(x-1) 1 can you fill in the other stuff?

OpenStudy (anonymous):

Okay . 12 seats . f(x)=12. f(x)=3(12-1)

OpenStudy (phi):

row # # of seats f(x) s(x)=3(x-1) 1 12 12 what goes in the last column? figure out s(1) which is 3(1-1)

OpenStudy (anonymous):

Thats 2

OpenStudy (phi):

3*(1-1) is not 2

OpenStudy (phi):

order of operations, first do stuff inside parens. what is 1 - 1 ?

OpenStudy (anonymous):

I was thinking to distribute

OpenStudy (phi):

if you distribute , 3*(1-1) becomes 3*1 - 3*1 order of operations: do the multiplies to get 3 - 3 now the subtraction

OpenStudy (anonymous):

Oh right.... darn...

OpenStudy (anonymous):

So we have 3. What next?

OpenStudy (phi):

3-3 is not 3

OpenStudy (anonymous):

That one was a typo, hahaa. Its 0

OpenStudy (phi):

ok, to be clearer, x is the row # x=row # # of seats f(x) s(x)=3(x-1) 1 12 12 0 2 now fill in the next row, for row # 2 (meaning x=2)

OpenStudy (anonymous):

15 15 3

OpenStudy (phi):

are you sure f(x)= 15 ?? they don't say that

OpenStudy (anonymous):

The say every row has 3 more seats than the one before it

OpenStudy (phi):

yes, so 15 is ok for the number of seats in row 2 but they also say f(x) = 12 (notice that no matter what x is , we will get 12) x=row # # of seats f(x) s(x)=3(x-1) 1 12 12 0 2 15 12 3 3 18 12 6 do you see how to figure out the # of seats in row 3 using f(x) and s(x) ?

OpenStudy (anonymous):

Oh okay

OpenStudy (phi):

do you see how to figure out the # of seats in row 3 using f(x) and s(x) ?

OpenStudy (phi):

in fact, do you see how to figure out the # of seats in row 1, or row 2 using f(x) and s(x)?

OpenStudy (anonymous):

Yes. There is one thing I cant understand. What does the outcome of the equation stand for?

OpenStudy (phi):

the outcome of what equation ?

OpenStudy (anonymous):

s(x)=3(x-1)

OpenStudy (phi):

There are 2 ways to answer that. First, look at the numbers in the table. What is s(x) telling us compared to the number of seats in the row ? or, try to interpret what they tell us The number of additional seats a row has.... that is a bit unclear, but based on the numbers, it looks like it means the number of seats more than 12 that we use in a row

OpenStudy (phi):

but, how do we figure out the number of seats in row 1, 2, or 3 using f(x) and s(x) ?

OpenStudy (anonymous):

Crap. Thanks for the help phi but ive got to go. Family emergency. I think ill be able to figure it out from here. Thanks!!!